17. The stationery and greeting cards department was given the following data to use in developing a merchandise plan for the spring season (3 months season): Planned sales for the season $120,000 April 30 inventory 80,000 Initial markup percent 53.5% % of Season’s BOM Stock Sales Markdowns Month Sales Ratio % February 42% 2.9 12% March 33% 3.4 15% April 25% 3.5 17% Calculate: (a) Monthly planned sales (b) BOM stocks (c) Monthly markdowns $ (d) Season total markdown $ and Season Markdown % (e) Monthly and total season planned purchases at retail (f) Monthly and total season planned purchases at cost (g) Planned stock turnover rate for the season
Answer #1

Answer-(a):

Answer-(b):

Average inventory = $80,000

Average sales = $120,000 / 3 = $40,000

Basic stock = Average inventory - Average sales = $80,000 - 40,000 = 40,000

Answer-(c):

Monthly markdown = Sales * Markdown %

Answer-(d):

Season's total markdown = $17,088

Seasons markdown % = Total markdown / Season's sales * 100 = $17,088 / 120,000*100 = 14.24%